The length of a rectangular playing field is 5 ft less than twice its width.
If the perimeter of the playing field is 230 ft, find the length and width of the field.

Answer by checkley75(3666)

L=2W-5
PERIMETER=2L+2W
230=2(2W-5)+2W
230=4W-10+2W
6W=230+10
6W=240
W=240/6
W=40 ANSWER FOR THE WIDTH.
L=2*40-5
L=80-5
L=75 ANSWER FOR THE LENGTH.
230=2*40+2*75
230=80+150
230=230